# Reinforcement Learning for LLM Reasoning Under Memory Constraints

### v1

@@ Abstract
We explore reinforcement learning (RL) techniques to enhance reasoning within targeted problem spaces in large language models (LLMs) under memory and compute constraints. Our focus is on critic-free methods compatible with LoRA fine-tuning on a single 40GB GPU, a common limitation in academic settings. We introduce S-GRPO, a memory-efficient variant of Group Relative Policy Optimization, and T-SPMO, a token-level prefix matching strategy for fine-grained credit assignment. Despite limited resources, when used to fine-tune Qwen2-1.5B both methods significantly improve SVAMP benchmark accuracy from 46% to above 70% using LoRA training. T-SPMO also excels in multi-digit multiplication tasks, underscoring the potential of RL fine-tuning under hardware constraints. Additionally, we find that our full-token GRPO baseline under LoRA fine-tuning did not improve model performance (compared to base model) on either task, suggesting that our memory-efficient methods may act as a form of regularization that stabilizes training when only a small subset of parameters are updated.
I Introduction
@@
Large Language Models (LLMs) have demonstrated substantial improvements in mathematical and structured problem-solving through reinforcement learning (RL) fine-tuning on specific tasks [1]. However, standard RL approaches such as Proximal Policy Optimization (PPO) [2] and Group Relative Policy Optimization (GRPO) [3] can be impractical for researchers operating under limited computational budgets, for the following reasons: (1) both methods compute loss over the entire output trajectory, (2) they are typically designed for full model fine-tuning, and (3) PPO additionally requires a critic network, further increasing memory demands.
Moreover, supervised fine-tuning approaches rely on high-quality, human-annotated chain-of-thought data, which can be expensive or infeasible to collect at scale for diverse problem domains.
@@
In this work, we explore the question: Can RL fine-tuning improve reasoning performance on targeted tasks using only parameter-efficient updates and a single 40GB GPU? To this end, we develop and evaluate two critic-free methods tailored for constrained memory and computational environments. First, we introduce Stochastic-GRPO (S-GRPO), a lightweight variant of GRPO that samples tokens from output trajectories to contribute to the loss. Second, we propose Token-Specific Prefix Matching Optimization (T-SPMO), which assigns credit at a token granularity.
@@
To evaluate these methods, we fine-tune the open-weight Qwen2-1.5B model [4] using LoRA [5] on two reasoning benchmarks: the SVAMP dataset [6] and a multi-digit arithmetic task. Despite the constrained training setup, both S-GRPO and T-SPMO raise SVAMP test accuracy from 46% to above 70%. On the multiplication benchmark, S-GRPO improves accuracy from 3.9% to 22.9%, while T-SPMO dramatically outperforms both baselines, reaching 70%. To ground our results, we also establish a full-token GRPO baseline under LoRA, which surprisingly did not lead to meaningful jumps in performance from the base model. These results suggest that carefully designed, memory-efficient RL algorithms can significantly improve LLM reasoning under compute constraints—even outperforming exhaustive full-token optimization strategies.
Our contributions are threefold:
We present S-GRPO, a lightweight variant of GRPO with minimal alterations.
We propose T-SPMO, a novel token-level RL method with fine-grained credit assignment tailored for reasoning tasks.
We demonstrate that both methods are effective when applied with LoRA-based fine-tuning under tight GPU wall-clock time and memory budgets, making RL for reasoning accessible to a broader research community.
II Related Work
II-A RL for Large Language Models
@@
RL has been widely used to align language model outputs with user preferences or task-specific objectives. Proximal Policy Optimization (PPO) [2] is a commonly used algorithm in RL from human feedback (RLHF), most notably applied in training models like InstructGPT [7]. However, PPO requires training a separate value network (critic), which significantly increases memory and compute demands.
@@
Group Relative Policy Optimization (GRPO) [3] was introduced to remove the need for a critic by computing group-level statistics over generated samples. Although effective, GRPO still assumes access to full-model fine-tuning and typically requires the storage of full trajectories for reward normalization, limiting its applicability in low-resource settings.
II-B Memory-Efficient Fine-Tuning
@@
Parameter-efficient fine-tuning methods such as Low-Rank Adaptation (LoRA) [5] have enabled researchers to adapt large models on modest hardware by injecting trainable rank-decomposed matrices into frozen weights. These techniques allow significant reductions in memory usage and training time without substantially sacrificing downstream performance. Our work builds on this line of research by evaluating RL algorithms in the context of LoRA-based fine-tuning, rather than full backpropagation over all model weights.
II-C Reasoning Benchmarks
@@
Benchmarks such as GSM8K [8], SVAMP [6], and ASDiv [9] have been widely adopted to measure mathematical and symbolic reasoning in language models. SVAMP in particular requires verbal reasoning over numerical quantities and has been shown to be out-of-distribution (OOD) for many LLMs (such as by presenting irrelevant information in questions) [6].
III Methods
III-A Problem Setup
@@
We focus on enhancing the reasoning capabilities of large language models (LLMs) using RL under constrained computational resources. The task is framed as conditional generation with reward feedback, where the model generates a sequence of tokens $Y=(y_{1},y_{2},\dots,y_{T})$ in response to a prompt $X$, and receives a scalar reward signal $r(Y)\in\mathbb{R}$ based on the correctness or quality of the generated sequence.
@@
$\mathcal{J}(\theta)=\mathbb{E}_{Y\sim\pi_{\theta}}\left[r(Y)\right]$ (1)
Traditional policy gradient methods often rely on estimating the advantage function using a separate value network, which can be computationally intensive and memory-demanding. To address these challenges, the recently popularized critic-free RL algorithm Group Relative Policy Optimization (GRPO) estimates advantages by comparing the rewards of multiple responses generated for the same prompt [3].
@@
[…]
$\quad\min\Biggl{[}\frac{\pi_{\theta}\!\bigl{(}y_{t}^{(i)}\!\mid\!x,y_{<t}^{(i)}\bigr{)}}{\pi_{\theta_{\text{old}}}\!\bigl{(}y_{t}^{(i)}\!\mid\!x,y_{<t}^{(i)}\bigr{)}}\,\hat{A}_{i,t},$
@@
$\mathrm{D}_{\mathrm{KL}}\left[\pi_{\theta}\,\|\,\pi_{\text{ref}}\right]=\frac{\pi_{\text{ref}}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})}{\pi_{\theta}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})}-\log\frac{\pi_{\text{ref}}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})}{\pi_{\theta}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})}-1$ (4)
To further accommodate limited computational resources, we employ Low-Rank Adaptation (LoRA) [5] for parameter-efficient fine-tuning. LoRA allows us to fine-tune only a small subset of the model’s parameters, significantly reducing memory requirements.
@@
  • SVAMP, a benchmark of verbal arithmetic problems requiring symbolic and numerical reasoning [6].
@@
In the following sections, we describe two methods—S-GRPO and T-SPMO—that implement RL updates without requiring full-sequence baselines or critic networks, making them well-suited to low-memory training environments. We then discuss the implementation of our baseline GRPO, which sacrifices wall clock time for memory efficiency using gradient accumulation.
III-B S-GRPO: Stochastic Group Relative Policy Optimization
@@
We also set the number of updates per problem (denoted as $\mu$ in the original GRPO paper [3]) to 1, aligning with the default setting in lightweight RL libraries such as TRL. This choice both simplifies optimization by eliminating the need for PPO-style clipped objectives and reduces computational overhead, consistent with our focus on efficient fine-tuning. Also in alignment of the default setting in TRL, we do not update $\pi_{\text{ref}}$ to $\pi_{\theta}$.
@@
$\mathcal{J}_{\text{SGRPO}}(\theta)$ $=$ $\frac{1}{|\mathcal{G}|}\sum_{i=1}^{\mathcal{G}}\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}\frac{\pi_{\theta}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})}{\pi_{\theta}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})\big{|}_{\text{no grad}}}\hat{A}_{i,t}$ (5)
${}-\,\beta\,D_{\text{KL}}\!\bigl{[}\pi_{\theta}\;\|\;\pi_{\text{ref}}\bigr{]}$
@@
Algorithm 1 presents the corresponding pseudocode.
Algorithm 1 Token Selection for Loss Calculation
1:Token indices $t=1,2,\ldots,T$
2:Cutoff index $\alpha$
3:Maximum tokens $k$
4:Sampling probability $P$
5:Initialize $\mathcal{T}_{i}\leftarrow\emptyset$
6:for each token $t$ from $1$ to $T$ do
7:     if $t<\alpha$ then
8:         Include token: $\mathcal{T}_{i}\leftarrow\mathcal{T}_{i}\cup\{t\}$
9:     else if $|\mathcal{T}_{i}|<k$ then
10:         Sample $s\sim\text{Bernoulli}(P)$
11:         if $s=1$ then
12:              Include token: $\mathcal{T}_{i}\leftarrow\mathcal{T}_{i}\cup\{t\}$
13:         end if
14:     else
15:         break
16:     end if
17:end for
18:return $\mathcal{T}_{i}$
This hybrid rule ensures early tokens—which often contain key semantic steps—are reliably updated, while sampling later tokens only when there is enough space. This strikes a balance between learning stability and memory efficiency. This stochastic sampling mechanism allows S-GRPO to scale to larger batches without retaining full trajectories.
Note that in GRPO, all tokens $t$ are guaranteed to be in $\mathcal{T}_{i}$ [3].
III-C T-SPMO: Token-Specific Prefix Matching Optimization
@@
We approximate these expectations using empirical averages over the sampled completions. Let $\mathcal{G}$ be the set of all sampled completions for a given prompt. We define:
$\mathbb{E}[R\mid p]$ $=\frac{1}{|\mathcal{G}_{p}|}\sum_{y\in\mathcal{G}_{p}}R(y)$ (8)
$\mathbb{E}[R\mid p\circ v]$ $=\frac{1}{|\mathcal{G}_{p\circ v}|}\sum_{y\in\mathcal{G}_{p\circ v}}R(y)$ (9)
@@
$\mathcal{J}_{\text{TSPMO}}(\theta)=\frac{1}{|\mathcal{U}|}\sum_{(p,v)\in\mathcal{U}}\pi_{\theta}(v\mid p)\cdot A(v\mid p)-\lambda\sum_{W\in\mathcal{W}_{\text{LoRA}}}\|W\|_{2}^{2}$ (10)
@@
III-D Full GRPO Baseline
To isolate the impacts of sparse‐token sampling, we implement the original GRPO objective (Eq. 3). We again set the number of iterations per sample to 1. The full GRPO objective is thus equivalent to the S-GRPO objective (Eq. 5), except all tokens are included in $\mathcal{T}_{i}$. We adapt GRPO to our memory constraints using gradient accumulation: We generate $\mathcal{G}$ without gradient enabled, then for each completion in $\mathcal{G}$ we run a forward and backwards pass. See Algorithm 2 for details.
Algorithm 2 Memory Efficient GRPO Training (Per-Completion Backward)
1:Prompt $x$, number of completions $G$, maximum completion length $T$, optimizer
2:Sample $G$ completions $\{y^{(i)}\}_{i=1}^{G}$ from policy $\pi_{\theta}$ without gradients
3:Compute reward $r(y^{(i)})$ for each completion
4:Calculate advantages $\hat{A}_{i}$
5:optimizer.zero_grad()
6:for each sampled completion $y^{(i)}$ do
7:     Build input by concatenating $x$ and $y^{(i)}$
8:     Forward policy $\pi_{\theta}$ and reference $\pi_{\text{ref}}$ over full input
9:     Compute GRPO loss using logits of generated tokens, $\hat{A}_{i}$, and KL-divergence
10:     Divide loss by G
11:     loss.backward()
12:end for
13:optimizer.step()
IV Experiments and Results
IV-A Experiment Setup
@@ IV-B Results
TABLE I: SVAMP benchmark test-set (n=300) results showing accuracy and approximate wall-clock training time for each method.
Table I shows both S-GRPO and T-SPMO substantially improved performance on SVAMP over the base model. We also find that GRPO did not meaningfully enhance the model’s performance compared to the base model.
@@
TABLE II: 3 digit x 3 digit multiplication test-set (n=3000) results showing accuracy and approximate wall-clock time for each method.
As demonstrated in Table II, the base Qwen2-1.5B model struggled significantly on 3 digit x 3 digit multiplication. Our baseline GRPO also did not improve performance. We see noticeable improvement after training with S-GRPO, but an early plateau in training resulted in low albeit improved test set accuracy. T-SPMO on the other hand excelled on this benchmark, significantly outperforming the baseline and S-GRPO.
@@ IV-C Ablations
($\alpha$, $k$) SVAMP Accuracy (%) Multiplication Accuracy (%)
@@
TABLE III: S-GRPO ablations on $\alpha$ and $k$ values. Test set accuracies on SVAMP (n=300) and multiplication (n=3000).
@@
Effective Batch Size SVAMP Acc (%) Multiplication Acc (%)
@@
TABLE IV: S-GRPO ablations on effective batch size. Test set accuracies on SVAMP and multiplication benchmarks.
@@
($|\mathcal{G}|$, $C_{\text{Success}}$, $C_{\text{Failure}}$) SVAMP Acc (%) Multiplication Acc (%)
@@
TABLE V: T-SPMO ablations on group size and replay strategy. Test set accuracies on SVAMP and multiplication benchmarks.
@@
IV-D Discussion
We hypothesize that T-SPMO outperforms S-GRPO on the multi-digit multiplication benchmark due to differences in how each method handles token-level credit assignment and partial correctness. While both encourage chain-of-thought reasoning, S-GRPO applies group-level updates across sampled tokens, making it sensitive to isolated errors: a single mistake can disproportionately penalize an otherwise accurate sequence. Our observations suggest that this broad attribution can cause useful subskills to be forgotten, destabilizing intermediate token accuracy.
T-SPMO addresses this by computing advantages for specific (prefix, token) pairs, allowing correct decisions to be reinforced even when later errors occur. This finer-grained credit assignment likely underpins the substantial token-level and final-answer accuracy gains we observe for T-SPMO over S-GRPO on multiplication.

### v2

# Token-Efficient RL for LLM Reasoning

@@ Abstract
We propose reinforcement learning (RL) strategies tailored for reasoning in large language models (LLMs) under strict memory and compute limits, with a particular focus on compatibility with LoRA fine-tuning. Rather than relying on full-sequence updates or separate critic networks, we design critic-free methods that operate on a small, informative subset of output tokens to reduce memory usage and stabilize training. We introduce S-GRPO, a stochastic variant of Group Relative Policy Optimization, and T-SPMO, a token-level prefix matching approach for fine-grained credit assignment. Applied to Qwen2-1.5B, our methods raise accuracy on the SVAMP benchmark from 46% to over 70% and show strong performance on multi-digit multiplication. Surprisingly, full-token GRPO under LoRA fails to improve over the base model, suggesting that selective token-level optimization may act as an implicit regularizer in low-parameter training regimes.
Machine Learning, ICML
1 Introduction
@@
Large Language Models (LLMs) have demonstrated substantial improvements in mathematical and structured problem-solving through reinforcement learning (RL) fine-tuning on specific tasks (Jaech et al., 2024). However, standard RL approaches such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024) can be impractical for researchers operating under limited computational budgets, for the following reasons: (1) both methods compute loss over the entire output trajectory, (2) they are typically designed for full model fine-tuning, and (3) PPO additionally requires a critic network, further increasing memory demands.
@@
In this work, we design and evaluate RL algorithms that are explicitly tailored to low compute constraints: Partial parameter updates, limited GPU capacity, and the need to mitigate overfitting in low-capacity adaptation settings. First, we introduce Stochastic-GRPO (S-GRPO), a lightweight variant of GRPO that samples tokens from output trajectories to contribute to the loss. Second, we propose Token-Specific Prefix Matching Optimization (T-SPMO), which assigns credit at a token granularity.
@@
To evaluate these methods, we fine-tune the open-weight Qwen2-1.5B model (Yang et al., 2024) using LoRA (Hu et al., 2021) on two reasoning benchmarks: the SVAMP dataset (Patel et al., 2021) and a multi-digit multiplication task. Both S-GRPO and T-SPMO significantly raise test-set accuracy on both benchmarks. We also establish a full-token GRPO baseline under LoRA, which surprisingly did not improve performance from the base model. These results suggest that carefully designed, memory-efficient RL algorithms can significantly improve LLM reasoning under compute constraints—even outperforming exhaustive full-token optimization strategies.
2 Related Work
2.1 RL for Large Language Models
@@
RL has been widely used to align language model outputs with user preferences or task-specific objectives. Proximal Policy Optimization (PPO) (Schulman et al., 2017) is a commonly used algorithm in RL from human feedback (RLHF), most notably applied in training models like InstructGPT (Ouyang et al., 2022). However, PPO requires training a separate value network (critic), which significantly increases memory and compute demands.
@@
Group Relative Policy Optimization (GRPO) (Shao et al., 2024) was introduced to remove the need for a critic by computing group-level statistics over generated samples. Although effective, GRPO still assumes access to full-model fine-tuning and typically requires the storage of full trajectories for reward normalization, limiting its applicability in low-resource settings.
2.2 Memory-Efficient Fine-Tuning
@@
Parameter-efficient fine-tuning methods such as Low-Rank Adaptation (LoRA) (Hu et al., 2021) have enabled researchers to adapt large models on modest hardware by injecting trainable rank-decomposed matrices into frozen weights. These techniques allow significant reductions in memory usage and training time without substantially sacrificing downstream performance. Our work builds on this line of research by evaluating RL algorithms in the context of LoRA-based fine-tuning, rather than full backpropagation over all model weights.
2.3 Reasoning Benchmarks
@@
Benchmarks such as GSM8K (Cobbe et al., 2021) and SVAMP (Patel et al., 2021) have been widely adopted to measure mathematical and symbolic reasoning in language models. SVAMP in particular requires verbal reasoning over numerical quantities and has been shown to be out-of-distribution (OOD) for many LLMs (such as by presenting irrelevant information in questions) (Patel et al., 2021).
3 Methods
3.1 Problem Setup
We focus on enhancing the reasoning capabilities of large language models (LLMs) using RL under constrained computational resources.
@@ 3.1 Problem Setup
The task is framed as conditional generation with reward feedback, where the model generates a sequence of tokens $Y=(y_{1},y_{2},\dots,y_{T})$ in response to a prompt $X$, and receives a scalar reward signal $r(Y)\in\mathbb{R}$ based on the correctness or quality of the generated sequence.
@@
$\mathcal{J}(\theta)=\mathbb{E}_{Y\sim\pi_{\theta}}\left[r(Y)\right].$ (1)
Traditional policy gradient methods often rely on estimating the advantage function using a separate value network, which can be computationally intensive and memory-demanding. To address these challenges, the recently popularized critic-free RL algorithm Group Relative Policy Optimization (GRPO) estimates advantages by comparing the rewards of multiple responses generated for the same prompt (Shao et al., 2024).
@@
$\mathcal{J}(\theta)$ $=\frac{1}{|\mathcal{G}|}\sum_{i=1}^{|\mathcal{G}|}\frac{1}{|y^{(i)}|}\sum_{t=1}^{|y^{(i)}|}\Biggl{\{}\min\Biggl{[}\frac{\pi_{\theta}\!\bigl{(}y_{t}^{(i)}\!\mid\!x,y_{<t}^{(i)}\bigr{)}}{\pi_{\theta_{\text{old}}}\!\bigl{(}y_{t}^{(i)}\!\mid\!x,y_{<t}^{(i)}\bigr{)}}\,\hat{A}_{i,t},$ (3)
@@
$\mathrm{D}_{\mathrm{KL}}\left[\pi_{\theta}\,\|\,\pi_{\text{ref}}\right]$ $=\frac{\pi_{\text{ref}}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})}{\pi_{\theta}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})}$
$\quad-\log\frac{\pi_{\text{ref}}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})}{\pi_{\theta}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})}-1$ (4)
To further accommodate limited computational resources, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2021) for parameter-efficient fine-tuning. LoRA allows us to fine-tune only a small subset of the model’s parameters, significantly reducing memory requirements.
@@
  • SVAMP, a benchmark of verbal arithmetic problems requiring symbolic and numerical reasoning (Patel et al., 2021).
@@
3.2 S-GRPO: Stochastic Group Relative Policy Optimization
@@
We also set the number of updates per problem (denoted as $\mu$ in the original GRPO paper (Shao et al., 2024)) to 1, aligning with the default setting in lightweight RL libraries such as TRL. This choice both simplifies optimization by eliminating the need for PPO-style clipped objectives and reduces computational overhead, consistent with our focus on efficient fine-tuning. Also in alignment of the default setting in TRL, we do not update $\pi_{\text{ref}}$ to $\pi_{\theta}$.
@@
$\mathcal{J}_{\text{SGRPO}}(\theta)$ $=\frac{1}{|\mathcal{G}|}\sum_{i=1}^{\mathcal{G}}\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}\frac{\pi_{\theta}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})}{\pi_{\theta}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})\big{|}_{\text{no grad}}}\hat{A}_{i,t}$
$\quad-\,\beta\,D_{\text{KL}}\!\left[\pi_{\theta}\;\|\;\pi_{\text{ref}}\right]$ (5)
@@
This hybrid rule ensures early tokens—which often contain key semantic steps—are reliably updated, while sampling later tokens only when there is enough space. This strikes a balance between learning stability and memory efficiency.
Note that in GRPO, all tokens $t$ are guaranteed to be in $\mathcal{T}_{i}$ (Shao et al., 2024).
3.3 T-SPMO: Token-Specific Prefix Matching Optimization
@@
We approximate these expectations using empirical averages over the sampled completions. Let $\mathcal{G}$ be the set of all sampled completions for a given prompt. We define
$\mathbb{E}[R\mid p]$ $=\frac{1}{|\mathcal{G}_{p}|}\sum_{y\in\mathcal{G}_{p}}R(y),$ (8)
$\mathbb{E}[R\mid p\circ v]$ $=\frac{1}{|\mathcal{G}_{p\circ v}|}\sum_{y\in\mathcal{G}_{p\circ v}}R(y),$ (9)
@@
$\mathcal{J}_{\text{TSPMO}}(\theta)$ $=\frac{1}{|\mathcal{U}|}\sum_{(p,v)\in\mathcal{U}}\pi_{\theta}(v\mid p)\cdot A(v\mid p)$
$\quad-\lambda\sum_{W\in\mathcal{W}_{\text{LoRA}}}\|W\|_{2}^{2}$ (10)
@@
3.4 Full GRPO Baseline
To isolate the impacts of sparse‐token sampling, we implement the original GRPO objective (Eq. 3). We again set the number of iterations per sample to 1. The full GRPO objective is thus equivalent to the S-GRPO objective (Eq. 5), except all tokens are included in $\mathcal{T}_{i}$. We adapt GRPO to low-memory environments using gradient accumulation: We generate $\mathcal{G}$ without gradient enabled, calculate rewards, then for each completion in $\mathcal{G}$ we run a forward and backward pass.
4 Experiments and Results
4.1 Experiment Setup
@@ 4.2 Results
Table 1: SVAMP benchmark test-set (n=300) results showing accuracy for each method.
@@
Table 1 shows both S-GRPO and T-SPMO substantially improved performance on SVAMP over the base model. We also find that GRPO did not meaningfully enhance the model’s performance compared to the base model.
Table 2: 3 digit x 3 digit multiplication test-set (n=3000) results showing accuracy for each method.
@@
As demonstrated in Table 2, the base Qwen2-1.5B model struggled significantly on 3 digit x 3 digit multiplication. Our baseline GRPO also did not improve performance. We see noticeable improvement after training with S-GRPO, but an early plateau in training resulted in low albeit improved test set accuracy. T-SPMO on the other hand excelled on this benchmark, significantly outperforming the baseline and S-GRPO.
@@ 4.3 Ablations
Table 3: S-GRPO ablations on $\alpha$ and $k$ values. Test set accuracies on SVAMP (n=300) and multiplication (n=3000).
($\alpha$, $k$) SVAMP Acc (%) Mult Acc (%)
@@
Table 4: S-GRPO ablations on effective batch size. Test set accuracies on SVAMP and multiplication benchmarks.
Batch Size SVAMP Acc (%) Mult Acc (%)
@@
Table 5: T-SPMO ablations on group size and replay strategy. Test set accuracies on SVAMP and multiplication benchmarks.
($|\mathcal{G}|$, $C_{\text{Suc}}$, $C_{\text{Fail}}$) SVAMP Acc (%) Mult Acc (%)
@@
4.4 Discussion
We hypothesize that T-SPMO outperforms S-GRPO on the multi-digit multiplication benchmark due to differences in how each method handles token-level credit assignment and partial correctness. While both encourage chain-of-thought reasoning, S-GRPO applies group-level updates across sampled tokens, making it sensitive to isolated errors: a single mistake can disproportionately penalize an otherwise accurate sequence.
T-SPMO addresses this by computing advantages for specific (prefix, token) pairs, allowing correct decisions to be reinforced even when later errors occur.